# Learning nonnegative matrix factorizations from compressed data

### v1

@@ 3.3. One-sided compression: nonorthogonal sketching matrices
The orthogonality assumption on ${\mathbf{A}}$ can be relaxed to having approximately orthogonal rows, such as those of appropriately normalized random matrices. This case is more than a straightforward extension of Theorem 3.4 because of the following computational challenge: if the sketching matrix ${\mathbf{A}}$ does not have orthogonal rows, the orthogonal projection operator $P^{\perp}_{\mathbf{A}}$ does not have a nicely decomposable form ${\mathbf{A}}^{T}{\mathbf{A}}$. Theorem 3.10 below shows how to having projection matrices in the regularizer term.
Definition 3.8 (Approximately orthogonal matrices).
For a positive constant $\varepsilon<1$, we call a matrix ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ $\varepsilon$-approximately orthogonal if its singular values lie in the interval $[1-\varepsilon,1+\varepsilon]$.
@@
$(1-\varepsilon)\|P_{\mathbf{A}}{\mathbf{M}}\|_{F}\leq\|{\mathbf{A}}{\mathbf{M}}\|_{F}\leq(1+\varepsilon)\|P_{\mathbf{A}}{\mathbf{M}}\|_{F}.$ (3.10)
@@
For a positive semidefinite matrix ${\mathbf{Z}}$, let $\sqrt{{\mathbf{Z}}}$ denote the unique positive semidefinite matrix such that $(\sqrt{{\mathbf{Z}}})^{2}={\mathbf{Z}}$. Then, if ${\mathbf{A}}={\mathbf{U}}\boldsymbol{\Sigma}{\mathbf{V}}^{T}$ is a compact SVD decomposition of ${\mathbf{A}}$, $\sqrt{{\mathbf{A}}^{T}{\mathbf{A}}}={\mathbf{V}}\boldsymbol{\Sigma}{\mathbf{V}}^{T}$ and $P_{\mathbf{A}}={\mathbf{V}}{\mathbf{V}}^{T}$. This implies $\|P_{\mathbf{A}}{\mathbf{M}}\|_{F}=\|{\mathbf{V}}^{T}{\mathbf{M}}\|_{F},$ $\|{\mathbf{A}}{\mathbf{M}}\|_{F}=\|\sqrt{{\mathbf{A}}^{T}{\mathbf{A}}}{\mathbf{M}}\|_{F}=\|\boldsymbol{\Sigma}{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}$ and
$(1-\varepsilon)\|P_{\mathbf{A}}{\mathbf{M}}\|_{F}$ $\leq(1-\|{\mathbf{I}}-\boldsymbol{\Sigma}\|)\|{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}\leq\|{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}-\|({\mathbf{I}}-\boldsymbol{\Sigma}){\mathbf{V}}^{T}{\mathbf{M}}\|_{F}$
$\leq\|\boldsymbol{\Sigma}{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}\leq\|\boldsymbol{\Sigma}\|\|{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}\leq(1+\varepsilon)\|P_{\mathbf{A}}{\mathbf{M}}\|_{F}.$
@@
Theorem 3.10.
(Approximately orthogonal ${\mathbf{A}}$) Let ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ be any matrix and let ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ be $\varepsilon$-approximately orthogonal, with $\varepsilon\leq 0.5$. Let ${\mathbf{U}}_{0}\in{\mathbb{R}}_{+}^{m\times r}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}_{+}^{r\times n}$ give a solution to the original NMF problem (1.1) of rank $r$ and ${\mathbf{X}}_{0}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$. If $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}$ solve the following compressed NMF problem with the same rank $r$
$\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}=\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\left[\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}_{F}+\lambda\|{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}\right].$ (3.11)
@@
$\frac{\left\|{\mathbf{X}}-\left(1+\lambda\right)\tilde{{\mathbf{X}}}\right\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}\leq c\left[\frac{\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\frac{\|P^{\perp}_{\mathbf{A}}{\mathbf{X}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\varepsilon^{2}\right]$ (3.12)
@@
Approximate orthogonality in the form of (3.10) applied to the matrices ${\mathbf{M}}={\mathbf{X}}-\tilde{{\mathbf{X}}}$ and ${\mathbf{M}}={\mathbf{X}}-{\mathbf{Y}}$ allows to orthogonalize this inequality:
@@
where ${\mathbf{Q}}$ denotes the $k\times m$ matrix with orthogonal rows such that $P_{\mathbf{A}}={\mathbf{Q}}^{T}{\mathbf{Q}}$. Indeed, this implies $\|P_{\mathbf{A}}{\mathbf{M}}\|_{F}=\|{\mathbf{Q}}^{T}{\mathbf{Q}}{\mathbf{M}}\|_{F}=\|{\mathbf{Q}}{\mathbf{M}}\|_{F}$ for any ${\mathbf{M}}\in{\mathbb{R}}^{m\times n}.$ So,
$\|{\mathbf{Q}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\delta\|\tilde{{\mathbf{X}}}\|^{2}_{F}\leq\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+\delta\|{\mathbf{Y}}\|^{2}_{F}+3\varepsilon^{2}\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}.$ (3.13)
@@
We will further rearrange the optimality condition using the following identity based on completion of the square on both sides of (3.13): for any matrices ${\mathbf{M}}_{1}$, ${\mathbf{M}}_{2}$ of appropriate size,
@@
Using this identity for ${\mathbf{M}}_{1}={\mathbf{Q}}{\mathbf{X}}$ and ${\mathbf{M}}_{2}={\mathbf{Q}}\tilde{{\mathbf{X}}}$ on the left and ${\mathbf{M}}_{2}={\mathbf{Q}}{\mathbf{Y}}$ on the right of (3.13), we obtain
$\frac{\delta}{1+\delta}\|{\mathbf{Q}}{\mathbf{X}}\|_{F}^{2}+\frac{1}{1+\delta}\|{\mathbf{Q}}({\mathbf{X}}-(1+\delta)\tilde{{\mathbf{X}}})\|^{2}_{F}+\delta\|P_{\mathbf{A}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}$
$\leq\frac{\delta}{1+\delta}\|{\mathbf{Q}}{\mathbf{X}}\|_{F}^{2}+\frac{1}{1+\delta}\|{\mathbf{Q}}({\mathbf{X}}-(1+\delta){\mathbf{Y}})\|^{2}_{F}+\delta\|P_{\mathbf{A}}^{\perp}{\mathbf{Y}}\|^{2}_{F}+3\varepsilon^{2}\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}.$
Cancelling common terms, letting ${\mathbf{Y}}:={\mathbf{X}}_{0}/(1+\delta)$, we have
$\frac{1}{1+\delta}\|{\mathbf{Q}}({\mathbf{X}}$ $-(1+\delta)\tilde{{\mathbf{X}}})\|^{2}_{F}+\frac{\delta}{(1+\delta)^{2}}\|P_{\mathbf{A}}^{\perp}(1+\delta)\tilde{{\mathbf{X}}}\|^{2}_{F}$
$\leq\frac{1}{1+\delta}\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+\delta\|P_{\mathbf{A}}^{\perp}\frac{{\mathbf{X}}_{0}}{1+\delta}\|^{2}_{F}+3\varepsilon^{2}\|{\mathbf{Q}}({\mathbf{X}}-\frac{{\mathbf{X}}_{0}}{1+\delta})\|^{2}_{F}=:{\mathbf{W}}.$
@@
To estimate the loss on the uncompressed problem, we use (3.7) with the matrices ${\mathbf{X}}$, $(1+\delta)\tilde{{\mathbf{X}}}$ and ${\mathbf{Q}}$ to get
$\|{\mathbf{X}}-(1+\delta)\tilde{{\mathbf{X}}}\|^{2}_{F}\leq\|{\mathbf{Q}}({\mathbf{X}}-(1+\delta)\tilde{{\mathbf{X}}})\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}(1+\delta)\tilde{{\mathbf{X}}}\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\quad\quad\quad\quad\quad\quad\leq\frac{2(1+\delta)^{2}}{\delta}{\mathbf{W}}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq\frac{2(1+\delta)}{\delta}\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}_{0}\|^{2}_{F}+\frac{6\varepsilon^{2}}{\delta}\|{\mathbf{Q}}((1+\delta){\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq\left(\frac{2+12\varepsilon^{2}}{\delta}+2\right)\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+4\|P_{\mathbf{A}}^{\perp}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+6\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}+12\varepsilon^{2}\delta\|{\mathbf{Q}}{\mathbf{X}}\|_{F}^{2}$
$\leq\left(4+\frac{5}{4\lambda}\right)\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+6\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}+12\frac{\varepsilon^{2}\lambda}{(1-\varepsilon)^{2}}\|{\mathbf{X}}\|_{F}^{2},$
using that $\delta=\frac{\lambda}{(1-\varepsilon)^{2}}$ and $\varepsilon\leq\frac{1}{2}$. ∎
Remark 3.11.
@@
We conclude the section with the discussion of Theorem 3.10 result. We note that
@@
  • Theorem 3.10 shows it is possible to regularize the compressed NMF problem without the projection operator and to find a $(1+\lambda)$-rescaled factors. Note that the rescaling does not affect the learned nonnegative low-rank structure.
  • The property (3.10) $\|P_{\mathbf{A}}{\mathbf{M}}\|^{2}_{F}\sim\|{\mathbf{A}}{\mathbf{M}}\|^{2}_{F}$ is significantly more relaxed than the standard geometry preservation properties of the form $\|{\mathbf{M}}\|^{2}_{F}\sim\|{\mathbf{A}}{\mathbf{M}}\|^{2}_{F}$, such as Johnson-Lindenstrauss, oblivious subspace embedding, or restricted isometry property. The latter won’t be satisfied for, e.g., random Gaussian matrix ${\mathbf{A}}$ and arbitrary nonnegative rank $r$ matrices ${\mathbf{M}}$ (as needed within Theorem 3.10), unless there is no compression and $k\geq m$.
@@
  • The approximate orthogonality property (3.10) is not hard to satisfy with generic random matrices. For example, an i.i.d. Gaussian matrix ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ with each entry having mean 00 and variance $\frac{1}{m}$ is $\varepsilon$-approximately orthogonal with probability $1-2\exp(-\varepsilon^{2}m/8)$ as soon as $k\leq m\varepsilon^{2}/4$ (by [41, Corollary 5.35]).
@@
  • While it is easy to guarantee approximate orthogonality with generic matrices ${\mathbf{A}}$ (not learned from ${\mathbf{X}}$), the term $P_{\mathbf{A}}^{\perp}{\mathbf{X}}$ is still the part of the error bound. So, data-oblivious one-sided compression in general is not expected to result in exact recovery even if data matrix ${\mathbf{X}}$ admits exact nonnegative factorization.

### v2

@@ 3.3. One-sided compression: nonorthogonal sketching matrices
The orthogonality assumption on $\mathbf{A}$} can be relaxed to having approximately orthogonal rows, such as those of appropriately normalized random matrices. This case is more than a straightforward extension of Theorem 2 because of the following computational challenge: if the sketching matrix $\mathbf{A}$} does not have orthogonal rows, the orthogonal projection operator $P^{\perp}_{{\mathbf{A}}^{T}}$ does not have a nicely decomposable form ${\mathbf{A}}^{T}{\mathbf{A}}$. Theorem 4 below shows how to having projection matrices in the regularizer term.
Definition 1 (Approximately orthogonal matrices).
For a positive constant $\varepsilon<1$, we call a matrix ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ $\varepsilon$-approximately orthogonal if its nonzero singular values lie in the interval $1-\varepsilon,1+\varepsilon]$.
@@
(11) $(1-\varepsilon)\|P_{{\mathbf{A}}^{T}}{\mathbf{M}}\|_{F}\leq\|{\mathbf{A}}{\mathbf{M}}\|_{F}\leq(1+\varepsilon)\|P_{{\mathbf{A}}^{T}}{\mathbf{M}}\|_{F}.$
@@
For a positive semidefinite matrix $\mathbf{Z}$}, let $\sqrt{{\mathbf{Z}}}$ denote the unique positive semidefinite matrix such that $(\sqrt{{\mathbf{Z}}})^{2}={\mathbf{Z}}$. Then, if ${\mathbf{A}}={\mathbf{U}}\boldsymbol{\Sigma}{\mathbf{V}}^{T}$ is a compact SVD decomposition of $\mathbf{A}$}, $\sqrt{{\mathbf{A}}^{T}{\mathbf{A}}}={\mathbf{V}}\boldsymbol{\Sigma}{\mathbf{V}}^{T}$ and $P_{{\mathbf{A}}^{T}}={\mathbf{V}}{\mathbf{V}}^{T}$. This implies $\|P_{{\mathbf{A}}^{T}}{\mathbf{M}}\|_{F}=\|{\mathbf{V}}^{T}{\mathbf{M}}\|_{F},$ $\|{\mathbf{A}}{\mathbf{M}}\|_{F}=\|\sqrt{{\mathbf{A}}^{T}{\mathbf{A}}}{\mathbf{M}}\|_{F}=\|\boldsymbol{\Sigma}{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}$ and
$(1-\varepsilon)\|P_{{\mathbf{A}}^{T}}{\mathbf{M}}\|_{F}$ $\leq(1-\|{\mathbf{I}}-\boldsymbol{\Sigma}\|)\|{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}\leq\|{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}-\|({\mathbf{I}}-\boldsymbol{\Sigma}){\mathbf{V}}^{T}{\mathbf{M}}\|_{F}$
$\leq\|\boldsymbol{\Sigma}{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}\leq\|\boldsymbol{\Sigma}\|\|{\mathbf{V}}^{T}{\mathbf{M}}\|_{F}\leq(1+\varepsilon)\|P_{{\mathbf{A}}^{T}}{\mathbf{M}}\|_{F}.$
@@
Theorem 4.
(Approximately orthogonal $\mathbf{A}$}) Let ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ be any matrix and let ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ be $\varepsilon$-approximately orthogonal, with $\varepsilon\leq 0.5$. Let ${\mathbf{U}}_{0}\in{\mathbb{R}}_{+}^{m\times r}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}_{+}^{r\times n}$ give a solution to the original NMF problem (1) of rank $r$ and ${\mathbf{X}}_{0}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$. If $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}$ solve the following compressed NMF problem with the same rank $r$
(12) $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}{\color[rgb]{0,0,0}\;\in\;}\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\left[\|{\mathbf{A}}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}_{F}+\lambda\|{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}\right],$
@@
(13) $\frac{\left\|{\mathbf{X}}-\left(1+\lambda\right)\tilde{{\mathbf{X}}}\right\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}\leq c\left[\frac{\|{\mathbf{X}}-{\color[rgb]{0,0,0}{\mathbf{X}}_{0}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\frac{\|P^{\perp}_{{\mathbf{A}}^{T}}{\mathbf{X}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\varepsilon^{2}\right]$
@@
Approximate orthogonality in the form of (11) applied to the matrices ${\mathbf{M}}={\mathbf{X}}-\tilde{{\mathbf{X}}}$ and ${\mathbf{M}}={\mathbf{X}}-{\mathbf{Y}}$ allows to orthogonalize this inequality:
@@
where $\mathbf{Q}$} denotes the $k\times m$ matrix with orthogonal rows such that $P_{{\mathbf{A}}^{T}}={\mathbf{Q}}^{T}{\mathbf{Q}}$. Indeed, this implies
(14) $\|P_{{\mathbf{A}}^{T}}{\mathbf{M}}\|_{F}=\|{\mathbf{Q}}^{T}{\mathbf{Q}}{\mathbf{M}}\|_{F}=\|{\mathbf{Q}}{\mathbf{M}}\|_{F}$
for any ${\mathbf{M}}\in{\mathbb{R}}^{m\times n}.$ So,
(15) $\|{\mathbf{Q}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\delta\|\tilde{{\mathbf{X}}}\|^{2}_{F}\leq\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}+\delta\|{\mathbf{Y}}\|^{2}_{F}+3\varepsilon^{2}\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}.$
@@
We will further rearrange the optimality condition using the following identity based on completion of the square on both sides of (15): for any matrices ${\mathbf{M}}_{1}$, ${\mathbf{M}}_{2}$ of appropriate size,
@@
Using this identity for ${\mathbf{M}}_{1}={\mathbf{Q}}{\mathbf{X}}$ for both sides and ${\mathbf{M}}_{2}={\mathbf{Q}}\tilde{{\mathbf{X}}}$ on the left and ${\mathbf{M}}_{2}={\mathbf{Q}}{\mathbf{Y}}$ on the right of (15), we obtain
$\frac{\delta}{1+\delta}\|{\mathbf{Q}}{\mathbf{X}}\|_{F}^{2}+\frac{1}{1+\delta}\|{\mathbf{Q}}({\mathbf{X}}-(1+\delta)\tilde{{\mathbf{X}}})\|^{2}_{F}+\delta\|P_{{\mathbf{A}}^{T}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}$
$\leq\frac{\delta}{1+\delta}\|{\mathbf{Q}}{\mathbf{X}}\|_{F}^{2}+\frac{1}{1+\delta}\|{\mathbf{Q}}({\mathbf{X}}-(1+\delta){\mathbf{Y}})\|^{2}_{F}+\delta\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{Y}}\|^{2}_{F}+3\varepsilon^{2}\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{Y}})\|^{2}_{F}.$
Canceling common terms, letting ${\mathbf{Y}}:={\mathbf{X}}_{0}/(1+\delta)$, we have
$\frac{1}{1+\delta}$ $\left\|{\mathbf{Q}}({\mathbf{X}}-(1+\delta)\tilde{{\mathbf{X}}})\right\|^{2}_{F}+\frac{\delta}{(1+\delta)^{2}}\left\|P_{{\mathbf{A}}^{T}}^{\perp}(1+\delta)\tilde{{\mathbf{X}}}\right\|^{2}_{F}$
$\leq\frac{1}{1+\delta}\left\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{X}}_{0})\right\|^{2}_{F}+\delta\left\|P_{{\mathbf{A}}^{T}}^{\perp}\frac{{\mathbf{X}}_{0}}{1+\delta}\right\|^{2}_{F}+3\varepsilon^{2}\left\|{\mathbf{Q}}({\mathbf{X}}-\frac{{\mathbf{X}}_{0}}{1+\delta})\right\|^{2}_{F}=:{\mathbf{W}}.$
@@
To estimate the loss on the uncompressed problem, we use (8) and (14) with the matrices $\mathbf{X}$}, $(1+\delta)\tilde{{\mathbf{X}}}$ and $\mathbf{Q}$} to get
$\|{\mathbf{X}}-(1+\delta)\tilde{{\mathbf{X}}}\|^{2}_{F}$ $\leq\|{\mathbf{Q}}({\mathbf{X}}-(1+\delta)\tilde{{\mathbf{X}}})\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}(1+\delta)\tilde{{\mathbf{X}}}\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq\frac{2(1+\delta)^{2}}{\delta}{\mathbf{W}}+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq\frac{2(1+\delta)}{\delta}\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+\frac{6\varepsilon^{2}}{\delta}\|{\mathbf{Q}}((1+\delta){\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}_{0}\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}.$
Now, since $\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}_{0}\|^{2}_{F}\leq 2\|P_{{\mathbf{A}}^{T}}^{\perp}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$, and then using $\delta=\lambda/(1-\varepsilon)^{2}$ and $\varepsilon\leq 0.5$, we can finally bound $\|{\mathbf{X}}-(1+\delta)\tilde{{\mathbf{X}}}\|^{2}_{F}$ by
$\left(\frac{2+12\varepsilon^{2}}{\delta}+2\right)$ $\|{\mathbf{Q}}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+4\|P_{{\mathbf{A}}^{T}}^{\perp}({\mathbf{X}}-{\mathbf{X}}_{0})\|^{2}_{F}+6\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}+12\varepsilon^{2}\delta\|{\mathbf{Q}}{\mathbf{X}}\|_{F}^{2}$
$\leq\left(4+\frac{5}{4\lambda}\right)\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+6\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}+12\frac{\varepsilon^{2}\lambda}{(1-\varepsilon)^{2}}\|{\mathbf{X}}\|_{F}^{2}.$
Remark 4.
@@
We conclude the section with the discussion of Theorem 4 result. We note that
  • Observe that the proofs of Theorem 4 and Theorem 2 above rely on the nonnegativity constraint only to ensure that the solution of the original NMF problem remains feasible for the compressed one.
@@
  • Theorem 4 shows it is possible to regularize the compressed NMF problem without the projection operator and to find a $(1+\lambda)$-rescaled factors. Note that the rescaling does not affect the learned nonnegative low-rank structure.
@@
  • The property (11) $\|P_{{\mathbf{A}}^{T}}{\mathbf{M}}\|^{2}_{F}\sim\|{\mathbf{A}}{\mathbf{M}}\|^{2}_{F}$ is very different from the standard geometry preservation properties of the form $\|{\mathbf{M}}\|^{2}_{F}\sim\|{\mathbf{A}}{\mathbf{M}}\|^{2}_{F}$, such as Johnson-Lindenstrauss, oblivious subspace embedding, or restricted isometry property (which can be shown to fail for, e.g., random Gaussian matrices ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ as long as $k<m$). In contrast, the approximate orthogonality property (11) is not hard to satisfy with generic random matrices. For example, an i.i.d. Gaussian matrix ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ with each entry having mean 0 and variance $\frac{1}{m}$ is $\varepsilon$-approximately orthogonal with probability $1-2\exp(-\varepsilon^{2}m/8)$ as soon as $k\geq m\varepsilon^{2}/4$ (by [44, Corollary 5.35]).
@@
  • While it is easy to guarantee approximate orthogonality with generic matrices $\mathbf{A}$} (not learned from $\mathbf{X}$}), the term $P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}$ is still the part of the error bound. So, data-oblivious one-sided compression in general is not expected to result in exact recovery even if data matrix $\mathbf{X}$} admits exact nonnegative factorization.